Premiums. Irwin Music Shop gives its customers coupons redeemable for a poster plus a Bo Diddley CD. One coupon is issued for each dollar of sales. On the surrender of 100 coupons and $6.00 cash, the poster and CD are given to the customer. It is estimated that 80% of the coupons will be presented for redemption. Sales for the first period were $700,000, and the coupons redeemed totaled 420,000. Sales for the second period were $840,000, and the coupons redeemed totaled 750,000. Irwin Music Shop bought 20,000 posters at $2.50/poster and 20000 CDs at 7.50 dollars per cd. Prepare journal entries for period 1 and period 2. Record entry for coupons redeemed and estimated liability
Answer #1

(1)

period 1 period 2
estimated liability for premiums 5600   debit
premium expense 16800 DEBIT 24400 debit [5600+45000-75000]
cash 25200 DEBIT 45000 debit
inventory of posters and CD            42000 CREDIT        75000 credit
number of coupons redeemed 420000 sales/100coupons = 4200 750000/100=7500
cash received 4200*6=25200 7500*6 = 45000
cost per coupon [2.5+7.5] 10$ 10$
inventory of CD and poster reduced

2.5*4200=10500

7.5*4200=31500

7500*10=75000
Premium expense will be difference between cost of CD AND POSTER and cash recovered from cusomer [ 10$ -6$] = 4$ 4200coupons redeemed *4$= 16800 7500*4=30000

(2)

1 2
premium expense 5600 debit 2480 debit
    estimated liability for premiums 5600 credit 2480 credit

estimated liability for coupons not redeemed =

[(700000*80%)-420000 coupons rdeemed]/100coupons *4$ cost of redemption

5600$
840000*80%= 672000 for period 2

coupons redeemed in period 2 - coupons of period 1 redeemed in period 2

=750000- [560000-420000]

=610000$ of sales were redeemed for period 2

coupons for period 2 not redeemed=

estimated coupon redemption - actual redemption = 672000-610000

=62000

liability = 62000/100 * 4$

=2480$